OpenStudy (anonymous):

Solve by Factoring: x2 - 2x + 1 = 0 What are the solutions for x? x = 2, x = 1 x = 1, x = -1 x = 1 none of the above

OpenStudy (anonymous):

breezy - do you know how to factor? you just need to find 2 numbers that add up AND multiply to value of 1?

OpenStudy (anonymous):

is it x=1?

OpenStudy (anonymous):

Yes its x=1

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

I would like you to factor this out for me. I know the answer but thought you would like to learn to do this

OpenStudy (anonymous):

ok can i have an example first

OpenStudy (anonymous):

Yep, It doesn't really help JUST getting answers, because like my math teacher used to always say, it always seems to build up not knowing how to do one things collides into not knowing how to do another. It's a nightmare!

OpenStudy (anonymous):

Most def Alf - that is why I did not just give her the answer :)

OpenStudy (anonymous):

Breezy - factor the problem you gave x2 - 2x + 1 = 0

OpenStudy (anonymous):

1x1 = 1 1+0 = 1

OpenStudy (anonymous):

remeber they must be the same two numbers. sorry the adding must total -2 (sorry) you are correct -1 * -1 = 1 and -1 + -1 = -2 (x -1 ) (x -1 )